m varies directly as n. If n= 2/3, then m= 1/4.
Answer by Cromlix(4381)

Hi there,
If m varies directly as n
m = kn
Where k = constant of proportionality
1/4 = k2/3
k = (1/4)/(2/3)
k = 3/8
Formula
m = 3/8n